OpenStudy (saaarrrahhhh):

A motorboat travels 9 miles downstream (with the current) in 30 minutes. The return trip upstream (against the current) takes 90 minutes. Which system of equations can be used to find x, the speed of the boat in miles per hour, and y, the speed of the current in miles per hour? Recall the formula d = rt. 9 = 0.5(x – y) 9 = 1.5(x + y) 9 = 1.5(x – y) 9 = 0.5(x + y) 0.5 = 9(x – y) 1.5 = 9(x + y) 1.5 = 9(x – y) 0.5 = 9(x + y)

OpenStudy (anonymous):

I think it may be b but i am not entirely sure but im in 6th grade so please dont judge

OpenStudy (saaarrrahhhh):

anybody? please?

OpenStudy (bibby):

given, d=rt, we can discount c and d because our answer is of the form: 9=0.5*r 9=1.5*r given that downstream=b+c and that upstream = b-c (traveling against the current) go wi th b

OpenStudy (bibby):

our answer is in that form because the distance is 9 miles and the times are 0.5/1.5

OpenStudy (bibby):

aka half hour and 90 mins